# Hiding Latencies in Network-Based Image Loading for Deep Learning

### v1

@@ 1 Introduction
  • We conduct a comprehensive evaluation of our approach, demonstrating its implementation and benchmarking its performance through extensive experiments in local, medium and high-latency settings.
@@ 4.2 Comparative analysis at varying latencies
[…]
@@
The results of these experiments, presented below, compare the performance of the data loaders across varying latency conditions.
@@ 4.2.2 Training
Figure 3: Normalized reading throughput at varying latencies.
@@
[…]
@@ 4.3.1 Impact of out-of-order, incremental prefetching
Figure 4 highlights the significant variance in batch loading times between in-order and out-of-order prefetching. In the in-order case (top plot), when the prefetching queue is exhausted, the system experiences delays of up to several seconds while waiting for all transfers, including those over congested routes, to complete. This results in a cyclical pattern: once all transfers are completed, the queue is refilled, but it is quickly depleted again, triggering a new cycle. In contrast, the out-of-order approach (bottom plot) maintains a highly consistent batch loading time, staying always below 30 ms after the initial transient period.
@@
Figure 4: Comparison of batch loading times in high-latency networks: in-order vs. out-of-order prefetching.
@@ Acknowledgment
This work was supported the Italian Ministry of Health under the program grant H2Hub (Hybrid Hub: Cellular and computation models, micro- and nano-technologies for personalized innovative therapies, project code T4-AN-10).

### v2

@@ 1 Introduction
  • We conduct a comprehensive evaluation of our approach, demonstrating its implementation and benchmarking its performance through extensive experiments in local, medium and high-latency settings, comparing it against the state-of-the-art tools.
@@ 4.2 Comparative analysis at varying latencies
  • Medium: Data stored in Northern California, RTT $\simeq$ 20 ms.
  • High: Data stored in Stockholm, Sweden, RTT $\simeq$ 150 ms.
@@
To minimize AWS usage costs, each data loader was evaluated in a single test run for up to four epochs. The experimental results presented below compare the performance of the data loaders under varying latency conditions.
@@ 4.2.2 Training
Figure 3: Normalized reading throughput at varying latencies. The red, dashed line in (a) shows the available network bandwidth.
@@
DALI TFRecord 7381 $\pm$ 108
MosaicML SD 326 $\pm$ 14 308 $\pm$ 15 203 $\pm$ 812
tf.data service 437 $\pm$ 512 57 $\pm$ 123 12 $\pm$ 123
Cassandra-DALI 6066 $\pm$ 147 5957 $\pm$ 115 4081 $\pm$ 337
@@ 4.3.1 Impact of out-of-order, incremental prefetching
Figure 4 highlights the significant variance in batch loading times between in-order and out-of-order prefetching. In the in-order case (Figure 4(a)), when the prefetching queue is exhausted, the system experiences delays of up to several seconds while waiting for all transfers, including those over congested routes, to complete. This results in a cyclical pattern: once all transfers are completed, the queue is refilled, but it is quickly depleted again, triggering a new cycle. In contrast, the out-of-order approach (Figure 4(b)) maintains a highly consistent batch loading time, staying always below 30 ms after the initial transient period.
@@
(a) In-order prefetching.
(b) Out-of-order prefetching.
Figure 4: Comparison of batch loading times in high-latency networks.
@@ Acknowledgment
This work was supported by the Italian Ministry of Health under the program grant H2ub (Hybrid Hub: Cellular and computation models, micro- and nano-technologies for personalized innovative therapies, project code T4-AN-10) and by the Regione Autonoma della Sardegna, Sardegna Ricerche, under the program grant XDATA.